OpenStudy (anonymous):

can someone please help simplify this? 570km/5.21hrs it is probably one of these a. 120 km/hr to the south b.110 km/hr to the south c.121 km/hr to the south d. need more information

OpenStudy (shamim):

B

OpenStudy (lncognlto):

Can you use a calculator?

OpenStudy (lncognlto):

I mean are you allowed to use one?

OpenStudy (shamim):

570/5.21

OpenStudy (anonymous):

it didn't help must cuz it still has to be a fraction

OpenStudy (lncognlto):

Ah, well then. 570 / 5 and 21/100 =570 / 521/100 = 570 x 100/521 =...